# Beam Cross Sections Create Mixtures: Improving Feature Localization in Secondary Electron Imaging

### v1

@@ IV-B Analyses of Time-Resolved Measurements
We now turn to analyses of the time-resolved measurement (25). As discussed in Section III-A, we only observe an SE count when it is positive, resulting in observed ion count $\widetilde{M}$ with distribution given in (21) and zero-truncated SE count distribution $\widetilde{X}$. Recall from Section III-D that given $\widetilde{M}$ and $\{\widetilde{X}_{1},\widetilde{X}_{2},\ldots,\widetilde{X}_{\widetilde{M}}\}$, the detection times $\{\widetilde{T}_{1},\widetilde{T}_{2},\ldots,\widetilde{T}_{\widetilde{M}}\}$ contain no information about the $X$ distribution.
@@
$\mathcal{I}_{\mathrm{TRM}}(q\,;\,\eta_{1},\eta_{2},\lambda)$
$=\mathcal{I}_{\widetilde{M}}(q\,;\,\eta_{1},\eta_{2},\lambda)+\mathrm{E}\!\left[\,{\widetilde{M}\,;\,\eta_{1},\eta_{2},\lambda}\,\right]\mathcal{I}_{\widetilde{X}_{i}}(q\,;\,\eta_{1},\eta_{2}).$ (56)
Two of the three quantities here are very simple. For the first term, we have the FI about $\rho$ from (37) that can be reparametrized using (53). Since $\partialderivative{\rho}{q}=e^{-\eta_{2}}-e^{-\eta_{1}}$, we have
$\mathcal{I}_{\widetilde{M}}(q\,;\,\eta_{1},\eta_{2},\lambda)=\frac{\lambda(e^{-\eta_{2}}-e^{-\eta_{1}})^{2}}{\rho}.$ (57)
@@
Refer to caption Refer to caption Refer to caption
Figure 5: The information gain factor from time-resolved measurement is the ratio of the FI in TRM to the FI in Poisson-compounded measurement $Y$ at high dose $\lambda$. The gain $\beta_{\mathrm{Poisson}}$ for Poisson-distributed $X$ is given in (34), and the gain $\beta_{\mathrm{mixture}}$ for mixture-distributed $X$ is given in (65).
@@
$\mathcal{I}_{\mathrm{TRM}}(q\,;\,\eta_{1},\eta_{2},\lambda)=\lambda\sum_{\widetilde{x}=1}^{\infty}\frac{1}{\widetilde{x}!}\frac{\left(\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}-\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}+q\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}}.$ (64)
As anticipated, normalized FI $\mathcal{I}_{\mathrm{TRM}}(q\,;\,\eta_{1},\eta_{2},\lambda)/\lambda$ is identical to the low-$\lambda$ asymptote (48). This again theoretically supports the information gain from using TRM. In the case of Poisson-distributed SE counts, the multiplicative information gain has a simple and interpretable form given in (34). We can similarly define the multiplicative information gain
@@
$={\left(\!\frac{\eta+\eta^{2}}{(\eta_{2}-\eta_{1})^{2}}+q(1-q)\!\right)}{\sum_{\widetilde{x}=1}^{\infty}\frac{1}{\widetilde{x}!}\frac{\left(\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}-\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}+q\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}}}$ (65)
@@
for the mixture SE distribution. While it is obvious that $\beta_{\mathrm{Poisson}}$ in (34) is approximately linear in $\eta$, (65) is more difficult to understand. We examine it through comparisons to (34) for a mean-matched Poisson distribution of SE counts. First, we consider $\eta_{1}\approx\eta_{2}$, which makes the mixture close to a Poisson distribution with matching mean. By varying $\eta_{1}$ with $\eta_{2}=\eta_{1}+1$ and $q=\frac{1}{2}$, Fig. 5(a) shows that $\beta_{\mathrm{mixture}}\approx\beta_{\mathrm{Poisson}}$. Cases where the mixture is farther from approximately Poisson are less intuitive. Figure 5(b) varies $q$ with $\eta_{1}=2$ and $\eta_{2}=8$. Here the information gain from TRM $\beta_{\mathrm{mixture}}$ increases sharply for $q\rightarrow 0$ and $q\rightarrow 1$. Finally, Fig. 5(c) varies $\eta_{2}$ with $\eta_{1}=4$ and $q=\frac{1}{2}$. Here $\beta_{\mathrm{mixture}}\leq\beta_{\mathrm{Poisson}}$, with equality where the mixture reduces to a simple Poisson distribution at $\eta_{2}=4$.
To conclude our analyses that are focused on a single raster scan location, we would like to emphasize that accuracy of the SE distribution model is paramount. Since the underlying analyses do not include impacts from model mismatch, one should not conclude from Fig. 5 that choosing a Poisson model is superior to choosing a mixture model at points where the red $\beta_{\mathrm{Poisson}}$ curve is higher than the blue $\beta_{\mathrm{mixture}}$ curve. These plots are merely meant to provide some intuition about $\beta_{\mathrm{mixture}}$.
@@
In this paper, a Poisson distribution for $X$ often symbolizes naïvety because (a) it arises from the convolutional model that we argue to be inaccurate; and (b) it sometimes justifies the uncritical use of sample mean estimators. The variance reduction depicted in Fig. 3 is for exploiting or not exploiting the mixture model when the mixture model does hold in describing the data. We can extend that result from the idealization of a deterministic beam in Section II-E to the more realistic case with source shot noise by comparing the MSE of the baseline estimator $\widehat{\eta}_{\rm baseline}$ to the MSE of an efficient estimator, i.e., the reciprocal of $\mathcal{I}_{\mathrm{TRM}}(\eta\,;\,\eta_{1},\eta_{2},\lambda)$. The red and blue curves in Fig. 6 provide this comparison for mean dose $\lambda=1$. The black curve is the MSE lower bound (19) for deterministically having $m=1$ incident ion. In the $q\rightarrow 0$ and $q\rightarrow 1$ limits, the gap between the red and blue curves exactly matches $\beta_{\mathrm{Poisson}}$ from (34). The gap between the blue and black curves is explained precisely by the FI gap (50).
@@
Figure 6: MSEs in SE yield estimation when SE counts follow the mixture model (9), as a function of mixing parameter $q$, for $\eta_{1}=2$, $\eta_{2}=8$, and $\lambda=1$. The baseline estimator $\widehat{\eta}_{\rm baseline}$ (28) does not use TRM and hence also cannot exploit the mixture model. Performance of an efficient estimate from TRM is described by (64) (with appropriate scaling). For additional context, performance described by (17) represents efficient estimation and a deterministic incident beam with $m=1$ ion.
@@ V Edge Localization
Recall that (8) in Section II-D connects a mixture parameter $q$ to an edge location $\gamma$ in a two-valued sample as shown in Fig. 2(a). This type of two-valued sample is inspired by semiconductor inspection problems, where an edge could be between a metal nanostructure and a silicon substrate. Having established several results in Section IV on the estimation of $q$ in the presence of source and target shot noise, we are now ready to study the estimation of $\gamma$—theoretically, in Monte Carlo simulations, and through processing of experimental data from an HIM. The precision of such edge localization has implications for the feature size tolerances of nanofabricated electronic devices, with consequent effects on device performance and reliability [44]. The localization here is not limited to the grid scan locations and hence represents a type of super-resolution.
In Section V-A, we use (64) to find the Fisher information about the edge location $\gamma$ in data from a collection of raster scan locations. We also explore optimizing the beam width $\sigma_{\mathrm{b}}$ to minimize the localization error, in worst case among all potential edge locations. In Section V-B, we introduce three edge location estimates. The maximum likelihood estimator for $\gamma$ from time-resolved measurement data under the mixture distribution (9) for SE counts $X$ is our central contribution for edge localization, and we wish to compare it to alternatives that do not exploit our novel modeling. Current practice is represented by linear interpolation of the baseline estimate of SE yield at each scan location. We also introduces an MLE formulation for TRM data under the conventional convolutional model (12) for the effect of the beam cross-section; this is mismatched with respect to the more accurate mixture model. Comparisons based on Monte Carlo simulations appear in Section V-C. Finally, Section V-D shows that experimental data acquired at a nanoscale edge between gold and silicon on a helium ion microscope show the same trends in estimator performance as the Monte Carlo simulations.
@@ V-A Fisher Information about the Edge Location in TRM
$\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})$
@@
Consistent with many results in earlier sections, $\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})$ is proportional to the dose $\lambda$; thus, we further study this FI normalized by $\lambda$. The dependence on beam width $\sigma_{\mathrm{b}}$ and edge location $\gamma$ is more intricate. In Fig. 7(a), we plot $\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda$ for $\gamma$ varied continuously between 46 and 53 for four values of $\sigma_{\mathrm{b}}$. The normalized FI exhibits oscillations depending on proximity to the scan grid. The oscillations are larger when the beam is narrow (e.g., $\sigma_{\mathrm{b}}\in\{0.2,\,0.3\}$) because an edge location $\gamma$ far from the scan grid (e.g., at 49.5) will result in a very small fraction of incident ions striking near the edge, while the opposite will happen when $\gamma$ is near the scan grid. Increasing the beam width (e.g., $\sigma_{\mathrm{b}}\in\{0.4,0.5\}$) dampens the oscillations but also reduces the normalized FI.
(a) $\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda$ as a function of $\gamma$
(b) $\min_{\gamma}\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda$ as a function of $\sigma_{\mathrm{b}}$
Figure 7: Study of Fisher information about the edge location $\gamma$ in a full scan at $\{0,\,1,\,\ldots,\,99\}$. (a) Dependence of dose-normalized FI on $\gamma$ (plotted for $\eta_{1}=1$, $\eta_{2}=10$, and four values of beam width $\sigma_{\mathrm{b}}$). (b) Dependence of the minimum over $\gamma$ of the dose-normalized FI on $\sigma_{\mathrm{b}}$. The maximin optimal value of the beam width is $\sigma_{\mathrm{b}}^{*}=0.33$.
Since the edge location is not known a priori, it is naturally desirable to maximize the informativeness of the measurements for the worst-case edge location. Hence, we define an optimal beam width $\sigma_{\mathrm{b}}^{\star}$ through a maximin formulation:
$\sigma_{\mathrm{b}}^{\star}=\operatorname*{arg\,max}_{\sigma_{\mathrm{b}}}\,\min_{\gamma\in\Gamma}\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda,$ (69)
where $\Gamma$ is the set of possible edge locations. For $(\eta_{1},\eta_{2})=(1,10)$, Fig. 7(b) shows the minimum normalized FI for a range of beam widths, and it shows the maximum occurring at $\sigma_{\mathrm{b}}^{*}=0.33$. The dependence of $\sigma_{\mathrm{b}}^{*}$ on $(\eta_{1},\eta_{2})$ is mild.

### v2

@@ IV-B Analyses of Time-Resolved Measurements
We now turn to analyses of the time-resolved measurement (22). As discussed in Section III-A, we only observe an SE count when it is positive, resulting in observed ion count $\widetilde{M}$ with distribution given in (19) and zero-truncated SE count distribution $\widetilde{X}$. Recall from Section III-C that given $\widetilde{M}$ and $\{\widetilde{X}_{1},\widetilde{X}_{2},\ldots,\widetilde{X}_{\widetilde{M}}\}$, the detection times $\{\widetilde{T}_{1},\widetilde{T}_{2},\ldots,\widetilde{T}_{\widetilde{M}}\}$ contain no information about the $X$ distribution.
Refer to caption Refer to caption Refer to caption
Figure 5: The information gain factor $\beta_{\mathrm{mixture}}$ from time-resolved measurement is the ratio of the FI in TRM to the FI in Poisson-compounded measurement $Y$ at high dose $\lambda$ as given in (59). With $\eta_{1}\approx\eta_{2}$ as in (a), this gain factor is consistent with the $\beta_{\mathrm{Poisson}}$ factor for Poisson-distributed $X$ given in (29).
@@
$\mathcal{I}^{\mathrm{mix}}_{\mathrm{TRM}}(q\,;\,$ $\eta_{1},\eta_{2},\lambda)=\mathcal{I}^{\mathrm{mix}}_{\widetilde{M}}(q\,;\,\eta_{1},\eta_{2},\lambda)$
$+\ \mathrm{E}\!\left[\,{\widetilde{M}\,;\,\eta_{1},\eta_{2},\lambda}\,\right]\mathcal{I}^{\mathrm{mix}}_{\widetilde{X}_{i}}(q\,;\,\eta_{1},\eta_{2}).$ (50)
Two of the three quantities here are very simple. For the first term, we have the FI about $\rho$ from (32) that can be reparametrized using (47). Since ${{\partial\rho}/{\partial q}}=e^{-\eta_{2}}-e^{-\eta_{1}}$, we have
$\mathcal{I}^{\mathrm{mix}}_{\widetilde{M}}(q\,;\,\eta_{1},\eta_{2},\lambda)=\frac{\lambda(e^{-\eta_{2}}-e^{-\eta_{1}})^{2}}{\rho}.$ (51)
@@
$\mathcal{I}^{\mathrm{mix}}_{\mathrm{TRM}}(q\,;\,\eta_{1},\eta_{2},\lambda)=\lambda\sum_{\widetilde{x}=1}^{\infty}\frac{1}{\widetilde{x}!}\frac{\left(\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}-\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}+q\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}}.$ (58)
As anticipated, normalized FI $\mathcal{I}^{\mathrm{mix}}_{\mathrm{TRM}}(q\,;\,\eta_{1},\eta_{2},\lambda)/\lambda$ is identical to the low-$\lambda$ asymptote (43). This again theoretically supports the information gain from using TRM. In the case of Poisson-distributed SE counts, the multiplicative information gain has a simple and interpretable form given by $\beta_{\mathrm{Poisson}}$ in (29). We can similarly define a multiplicative information gain for the mixture SE distribution as the ratio between (43) and (44): [^4] While $\beta_{\mathrm{Poisson}}$ is a ratio of FIs about $\eta$, here we use FIs about $q$. Under our assumption that $\eta_{1}$ and $\eta_{2}$ are known, these differ from FIs about $\eta$ by a multiplicative factor that cancels in the ratio.
@@
$={\left(\!\frac{\eta+\eta^{2}}{(\eta_{2}-\eta_{1})^{2}}+q(1-q)\!\right)}$
$\qquad{\sum_{\widetilde{x}=1}^{\infty}\frac{1}{\widetilde{x}!}\frac{\left(\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}-\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{\widetilde{x}}e^{-\eta_{1}}+q\eta_{2}^{\widetilde{x}}e^{-\eta_{2}}}}.$ (59)
@@
While it is obvious that $\beta_{\mathrm{Poisson}}$ in (29) is approximately linear in $\eta$, (59) is more difficult to understand. We examine it through three plots in Fig. 5. Figure 5(a) shows $\beta_{\mathrm{mixture}}$ when varying $\eta_{1}$ with $\eta_{2}=\eta_{1}+1$ and $q=\frac{1}{2}$. The increase of $\beta_{\mathrm{mixture}}$ with increasing $\eta_{1}$ is quite similar to the dependence of $\beta_{\mathrm{Poisson}}$ on $\eta$. In fact, $\beta_{\mathrm{mixture}}$ approaches $\beta_{\mathrm{Poisson}}$ as $\eta_{1}$ and $\eta_{2}$ approach $\eta$. Cases where the mixture is farther from approximately Poisson are less intuitive. Figure 5(b) varies $q$ with $\eta_{1}=2$ and $\eta_{2}=8$. Here the information gain from TRM $\beta_{\mathrm{mixture}}$ increases sharply for $q\rightarrow 0$ and $q\rightarrow 1$. Finally, Fig. 5(c) varies $\eta_{2}$ with $\eta_{1}=4$ and $q=\frac{1}{2}$.
@@
In this paper, a Poisson distribution for $X$ often symbolizes naïvety because (a) it arises from the convolutional model that we argue to be inaccurate; and (b) it sometimes justifies the uncritical use of sample mean estimators. The variance reduction depicted in Fig. 3 is for exploiting or not exploiting the mixture model when the mixture model does hold in describing the data. We can extend that result from the idealization of a deterministic beam in Section II-E to the more realistic case with source shot noise by comparing the MSE of the baseline estimator $\widehat{\eta}_{\rm baseline}$ to the MSE of an efficient estimator, i.e., the reciprocal of $\mathcal{I}^{\mathrm{mix}}_{\mathrm{TRM}}(\eta\,;\,\eta_{1},\eta_{2},\lambda)$. The red and blue curves in Fig. 6 provide this comparison for mean dose $\lambda=1$. The black curve is the MSE lower bound implied by (17) for deterministically having $m=1$ incident ion. In the $q\rightarrow 0$ and $q\rightarrow 1$ limits, the gap between the red and blue curves exactly matches $\beta_{\mathrm{Poisson}}$ from (29). The gap between the blue and black curves is explained precisely by the FI gap (45).
@@
Figure 6: MSEs in SE yield estimation when SE counts follow the mixture model (7), as a function of mixing parameter $q$, for $\eta_{1}=2$, $\eta_{2}=8$, and $\lambda=1$. The baseline estimator $\widehat{\eta}_{\rm baseline}$ (25) does not use TRM and hence also cannot exploit the mixture model. Performance of an efficient estimate from TRM is described by (58) (with appropriate scaling). For additional context, performance described by (16) represents efficient estimation and a deterministic incident beam with $m=1$ ion. TRM enables substantial improvements as compared to the baseline estimator; a deterministic beam would enable further improvements, as described by the FI gap (45).
@@ V Edge Localization
Recall that (6) in Section II-D connects a mixture parameter $q$ to an edge location $\gamma$ in a two-valued sample as shown in Fig. 2(a). This type of two-valued sample is inspired by semiconductor inspection problems, where an edge could be between a metal nanostructure and a silicon substrate. Having established several results in Section IV on the estimation of $q$ in the presence of source and target shot noise, we are now ready to study the estimation of $\gamma$—theoretically, in Monte Carlo simulations, and through processing of experimental data from an HIM. The precision of such edge localization has implications for the feature size tolerances of nanofabricated electronic devices, with consequent effects on device performance and reliability [44]. The localization here is not limited to the grid scan locations and provides sub-pixel accuracy.
In Section V-A, we use (58) to find the Fisher information about the edge location $\gamma$ in data from a collection of raster scan locations. We also explore optimizing the beam width $\sigma_{\mathrm{b}}$ to minimize the localization error, in worst case among all potential edge locations. In Section V-B, we introduce four edge location estimates. The maximum likelihood estimator for $\gamma$ from time-resolved measurement data under the mixture distribution (7) for SE counts $X$ is our central contribution for edge localization, and we wish to compare it to several alternatives. Current practice is represented by linear interpolation of the baseline estimate of SE yield at each scan location. This is slightly improved by clipping as in Section II-E; since $\eta_{1}$ and $\eta_{2}$ known, SE yield estimates are clipped to $\eta_{1},\eta_{2}]$, eliminating the cases in which $\widehat{\eta}_{\rm baseline}$ is slightly less than $\eta_{1}$ or slightly greater than $\eta_{2}$. We also introduce an MLE formulation for TRM data under the Poisson distribution of the data, i.e., treating the effect of the beam cross-section as a convolution; this is mismatched with respect to the more accurate mixture model. Comparisons based on Monte Carlo simulations appear in Section V-C. Finally, Section V-D uses experimental data acquired at an edge between gold and silicon on a helium ion microscope to show the same trends in estimator performance as the Monte Carlo simulations.
@@ V-A Fisher Information about the Edge Location in TRM
$\mathcal{I}^{\mathrm{mix}}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})$
@@
Consistent with many results in earlier sections, $\mathcal{I}^{\mathrm{mix}}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})$ is proportional to the dose $\lambda$; thus, we further study this FI normalized by $\lambda$. The dependence on beam width $\sigma_{\mathrm{b}}$ and edge location $\gamma$ is more intricate. In Fig. 7(a), we plot $\mathcal{I}^{\mathrm{mix}}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda$ for $\gamma$ varied continuously between 46 and 53 for four values of $\sigma_{\mathrm{b}}$. The normalized FI exhibits oscillations depending on proximity to the scan grid. The oscillations are larger when the beam is narrow (e.g., $\sigma_{\mathrm{b}}\in\{0.2,\,0.3\}$) because an edge location $\gamma$ far from the scan grid (e.g., at 49.5) will result in a very small fraction of incident ions striking near the edge, while the opposite will happen when $\gamma$ is near the scan grid. Increasing the beam width (e.g., $\sigma_{\mathrm{b}}\in\{0.4,0.5\}$) dampens the oscillations but also reduces the normalized FI.
(a) $\mathcal{I}^{\mathrm{mix}}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda$ as a function of $\gamma$
(b) $\min_{\gamma}\mathcal{I}^{\mathrm{mix}}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda$ as a function of $\sigma_{\mathrm{b}}$
Figure 7: Study of Fisher information about the edge location $\gamma$ in a full scan at $\{0,\,1,\,\ldots,\,99\}$. (a) Dependence of dose-normalized FI on $\gamma$ (plotted for $\eta_{1}=1$, $\eta_{2}=10$, and four values of beam width $\sigma_{\mathrm{b}}$). FI is highest when the edge is near a scan grid point and lowest when the edge is near the midpoint between scan grid points. (b) Dependence of the minimum over $\gamma$ of the dose-normalized FI on $\sigma_{\mathrm{b}}$. The maximin optimal value of the beam width is $\sigma_{\mathrm{b}}^{\star}=0.33$.
Since the edge location is not known a priori, it is desirable to maximize the informativeness of the measurements for the worst-case edge location. Hence, we define an optimal beam width $\sigma_{\mathrm{b}}^{\star}$ through a maximin formulation:
$\sigma_{\mathrm{b}}^{\star}=\operatorname*{arg\,max}_{\sigma_{\mathrm{b}}}\,\min_{\gamma\in\Gamma}\mathcal{I}^{\mathrm{mix}}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})/\lambda,$ (63)
where $\Gamma$ is the set of possible edge locations. For $(\eta_{1},\eta_{2})=(1,10)$, Fig. 7(b) shows the minimum normalized FI for a range of beam widths, and it shows the maximum occurring at $\sigma_{\mathrm{b}}^{\star}=0.33$. This is not a universal constant. The dependence of $\sigma_{\mathrm{b}}^{\star}$ on $(\eta_{1},\eta_{2})$ is mild, as illustrated in Supplementary Note 3.